# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ Appendix A Lemmas
In this Section, we introduce the lemmas used in the proofs of the finite-sample propositions and corollaries. The first set of lemmas corresponds to Assumption 1, where the data is assumed to be independent and drawn from a sub-exponential distribution. The second set pertains to Assumption 2, which allows for weakly dependent data.
@@ A.1 Independent Data
In what follows, we introduce the lemmas that provide the concentration inequalities for the estimators and latent quantities involved in the nonparametric bounds. The generalized Bernstein inequality for sub-exponential variables is taken from Vershynin, (2018), the Hoeffding bound for Bernoulli random variables from Hoeffding, (1994), and the Dvoretzky–Kiefer–Wolfowitz inequality from Kosorok, (2008).
@@ A.2 Weakly Dependent Data
In this section, we present the definitions and lemmas relevant to weakly dependent data. The definition of the $\alpha$-mixing process, as well as the concentration inequalities used to derive nonparametric bounds for weakly dependent data drawn from sub-exponential distributions, are drawn from White, (2014), Merlevède et al., (2009), Rio, (2000) and Dedecker and Merlevède, (2007).
@@
Let $Y_{1},\cdots,Y_{n}$ be a strictly stationray real-valued sequence with common CDF. $F$ and assume the strong mixing coefficients $\upalpha(k)$ in (40) satisfies $\sum_{k\geq 1}\upalpha(k)^{1/2}<\infty$. Define the empirical CDF as per Eq. (36). Then for every $t>0$ and $n\geq 1$
@@
where $C_{\upalpha}=\sum\limits_{k\geq 1}\upalpha(k)^{1/2}<\infty$. In particular, if $\upalpha(k)=0$ for all $k\geq 1$ (the independent case) then $C_{\upalpha}=0$ and (47) reduces to (37).
@@
In particular, if $\upalpha(k)\equiv 0$ then $C_{\upalpha}=0$ and this reduces to the usual Azuma–Hoeffding bound $\Pr(\lvert\hat{p}_{n}-p\rvert\geq t)\leq 2e^{-nt^{2}/2}$.
@@ B.1 Proof of Lemma 5
From Section 2, Theorem 1 and Remark 1 of Dedecker and Merlevède, (2007), it is known that for any finite measure $\upmu$ and $p\geq 2$:
@@
Next we relate the $\tau$ coefficients to $\upalpha$-mixing. Inequality (4.1) in Section 4.1 of Dedecker and Merlevède, (2007) shows:
@@ B.2 Proof of Proposition 1
$\Re\in\bigg{[}\hat{\delta}_{1}\hat{p}_{1}+L^{(1)}\hat{p}_{0}-U^{(0)}\hat{p}_{1}-\hat{\delta}_{0}\hat{p}_{0},\;\hat{\delta}_{1}\hat{p}_{1}+U^{(1)}\hat{p}_{0}-L^{(0)}\hat{p}_{1}-\hat{\delta}_{0}\hat{p}_{0}\bigg{]}$ (48)
For $u=m_{0},\cdots,m_{1}$, to simultaneously obtain the $(1-\alpha_{u})\%$ confidence set for both the upper and lower bounds for the identification region (48), we must first find the confidence bounds with an appropriate significance level and combine them using Bonferroni inequalities, such that the combined confidence set has a $(1-\alpha_{u})\%$ coverage rate, or:
$\Pr\left([\mathcal{L}(\hat{\mathbf{\theta}}),\mathcal{U}(\hat{\mathbf{\theta}})]\subseteq\mathbb{I}(\Re)\right)\geq 1-\alpha_{u},\quad\text{with}\quad\alpha_{u}=\frac{\alpha}{\overline{\mathcal{M}}}.$ (49)
where $\hat{\mathbf{\theta}}=(\hat{\delta}_{1},\hat{\delta}_{0},\hat{p}_{1},\hat{p}_{0})^{\top}$ is a $4\times 1$ vector of estimators and $\mathbb{I}(\Re)$ is an interval $l(\Re),u(\Re)]$. In other words, we wish to obtain:
$\Pr\left(l(\Re)\leq\mathcal{L}(\hat{\theta})\right)\geq 1-\frac{\alpha_{u}}{2},\quad\text{and}\quad\Pr\left(u(\Re)\geq\mathcal{U}(\hat{\theta})\right)\geq 1-\frac{\alpha_{u}}{2}$ (50)
such that,
$\Pr\left(l(\Re)\leq\mathcal{L}(\hat{\theta})\cap u(\Re)\geq\mathcal{U}(\hat{\theta})\right)\geq 1-\alpha_{u}$ (51)
@@
$\begin{split}\Pr\left(l(\Re)\leq\mathcal{L}(\hat{\theta})\cap u(\Re)\geq\mathcal{U}(\hat{\theta})\right)&\geq 1-\Pr\left(l(\Re)>\mathcal{L}(\hat{\theta})\right)-\Pr\left(u(\Re)<\mathcal{U}(\hat{\theta})\right)\\ &\geq 1-\frac{\alpha_{u}}{2}-\frac{\alpha_{u}}{2}\\ &\geq 1-\alpha_{u}\end{split}$ (52)
where the significance levels are such that $1-\alpha_{u,1}-\alpha_{u,2}=1-\alpha_{u}/2-\alpha_{u}/2=1-0.05$. Thus,
@@
$\sqrt{N}(\hat{\theta}-\theta)\xrightarrow{D}N(0,\Omega_{\theta})$ (54)
@@
$\mathcal{L}(\hat{\theta})\approx\mathcal{L}(\theta)+\nabla\mathcal{L}(\theta)^{\top}(\hat{\theta}-\theta)$ (55)
$\mathcal{U}(\hat{\theta})\approx\mathcal{U}(\theta)+\nabla\mathcal{U}(\theta)^{\top}(\hat{\theta}-\theta)$ (56)
@@
$\nabla\mathcal{L}(\theta)=\left(\frac{\partial}{\partial\delta_{1}}\mathcal{L}(\theta),\frac{\partial}{\partial\delta_{0}}\mathcal{L}(\theta),\frac{\partial}{\partial p_{1}}\mathcal{L}(\theta),\frac{\partial}{\partial p_{0}}\mathcal{L}(\theta)\right)^{\top}$ (57)
with $\nabla\mathcal{U}(\theta)$ defined similarly. Therefore, taking the $\text{Var}(.)$ of both sides of the equations (55) and (56) yields:
$\begin{split}\text{Var}\left(\mathcal{L}(\hat{\theta})\right)&\approx\text{Var}\left(\mathcal{L}(\theta)+\nabla\mathcal{L}(\theta)^{\top}(\hat{\theta}-\theta)\right)\\ &=\text{Var}\left(\mathcal{L}(\theta)+\nabla\mathcal{L}(\theta)^{\top}\hat{\theta}-\nabla\mathcal{L}(\theta)^{\top}\theta)\right)\\ &=\text{Var}\left(\nabla\mathcal{L}(\theta)^{\top}\hat{\theta}\right)\\ &=\nabla\mathcal{L}(\theta)^{\top}cov\left(\hat{\theta}\right)\nabla\mathcal{L}(\theta)\\ &=\nabla\mathcal{L}(\theta)^{\top}\frac{\Omega_{0}}{N}\nabla\mathcal{L}(\theta)\end{split}$ (58)
@@
The covariance matrix of estimators $\hat{\mathbf{\theta}}$ is given explicitly by:
@@
Therefore,
$\sqrt{N}\left(\mathcal{L}(\hat{\theta})-\mathcal{L}(\theta)\right)$ $\xrightarrow{D}N\left(0,\nabla\mathcal{L}(\theta)^{\top}\Omega\nabla\mathcal{L}(\theta)\right)$ (62)
$\sqrt{N}\left(\mathcal{U}(\hat{\theta})-\mathcal{U}(\theta)\right)$ $\xrightarrow{D}N\left(0,\nabla\mathcal{U}(\theta)^{\top}\Omega\nabla\mathcal{U}(\theta)\right)$ (63)
B.3 Proof of Proposition 2
We wish to show how to obtain the coverage probability
$\Pr\left(\forall u\in\mathcal{M},\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha$ (64)
for $u=m_{0},\cdots,m_{1}$ and some arbitrary significance level $0<\alpha<1$ when assumption 1 is in place. To achieve this, we first need to consider the six “good” events:
$\mathcal{E}_{1}$ $:=\{\lvert\hat{\mu}_{1}-\mu_{1}\rvert\leq t_{1}\}$
$\mathcal{E}_{2}$ $:=\{\lvert\hat{\mu}_{0}-\mu_{0}\rvert\leq t_{2}\}$
$\mathcal{E}_{3}$ $:=\{\lvert\hat{p}_{1}-p_{1}\rvert\leq t_{3}\}$
$\mathcal{E}_{4}$ $:=\{\lvert\hat{p}_{0}-p_{0}\rvert\leq t_{4}\}$
$\mathcal{E}_{5}$ $:=\left\{\sup_{y}\bigg{\lvert}F_{N_{1}}^{(1)}(y)-F^{(1)}(y)\bigg{\rvert}\leq t_{5}\right\}$
$\mathcal{E}_{6}$ $:=\left\{\sup_{y}\bigg{\lvert}F_{N_{0}}^{(0)}(y)-F^{(0)}(y)\bigg{\rvert}\leq t_{6}\right\}$
Thus, showing $\Pr\left(\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha_{u}$ for $u=m_{0},\cdots,m_{1}$ is equivalent to showing that the intersection of the events, i.e., $\Pr(\bigcap\limits_{i=1}^{6}\mathcal{E}_{i})\geq 1-\alpha_{u}$. Using De Morgan’s law, it is clear that
@@ B.3 Proof of Proposition 2
$\geq 1-\sum\limits_{i=1}^{6}\Pr\left(\mathcal{E}_{i}^{c}\right)$ (68)
Hence, showing that the bound (64) holds is equivalent to ensuring that $\sum_{i=1}^{6}\Pr\left(\mathcal{E}_{i}^{c}\right)\leq\alpha_{u}$ for $u=m_{0},\cdots,m_{1}$. The only tools we need are the three inequalities in Lemmas 13.
@@
Let $N_{1}$ (resp. $N_{0}$) be the number of observations with $Z=1$ (resp. $Z=0$). Lemma 1 gives for any $t>0$
@@
(The first term is used when $t_{k}\leq M_{k}$ — the “quadratic” regime; otherwise the second, “linear”, term is smaller.) With this choice $2\exp[-\log(12/\alpha_{u})]=\alpha_{u}/6$, so $\Pr(\mathcal{E}_{1}^{c})=\Pr(\mathcal{E}_{2}^{c})=\alpha_{u}/6$.
@@
$\Pr\left(\lvert\hat{p}_{k}-p_{k}\rvert\geq t\right)\leq 2\exp[-2Nt^{2}],\qquad k=0,1.$
@@
$\Pr\left(\sup_{y}\lvert F^{(k)}_{n}(y)-F^{(k)}(y)\rvert>t\right)\leq 2\exp[-2N_{k}t^{2}],\qquad k=0,1.$
@@
This establishes (64).
B.4 Proof of Corollary 1
The argument follows Proposition 2 verbatim except that the empirical–CDF events are now one-sided because the lower support is the known constant $\lambda$:

### v2

@@ Appendix A Lemmas
In this section we first state a lemma that justifies the quantile “sandwich” bound for the empirical distribution function used in our nonparametric bounds for the latent conditional expectations. We then collect the lemmas used in the proofs of the finite-sample propositions and corollaries. The first set of lemmas in relation to the latter corresponds to Assumption 1, where the data is assumed to be independent and drawn from a sub-exponential distribution. The second set pertains to Assumption 2, which allows for weakly dependent data.
Lemma 1 (Quantile sandwich from uniform CDF control).
Let $F$ be a CDF and $\hat{F}_{N}$ its empirical CDF. On the event $\sup_{y}|\hat{F}_{N}(y)-F(y)|\leq\varepsilon$, for every $u\in[\varepsilon,1-\varepsilon]$,
$F^{-1}(u-\varepsilon)\ \leq\ \hat{F}_{N}^{-1}(u)\ \leq\ F^{-1}(u+\varepsilon),$
where $G^{-1}(u)=\inf\{y:\,G(y)\geq u\}$. In particular, $\hat{F}_{N}^{-1}(u)=Y_{(\lceil uN\rceil)}$ is the $\lceil uN\rceilth$ order statistic.
@@ A.1 Independent Data
In what follows, we introduce the lemmas that provide the concentration inequalities for the estimators and latent quantities involved in the nonparametric bounds. The generalized Bernstein inequality for sub-exponential variables is taken from vershynin2018high, the Hoeffding bound for Bernoulli random variables from hoeffding1994probability, and the Dvoretzky-Kiefer-Wolfowitz inequality from kosorok2008introduction.
@@ A.2 Weakly Dependent Data
In this section, we present the definitions and lemmas relevant to weakly dependent data. The definition of the $\alpha$-mixing process, as well as the concentration inequalities used to derive nonparametric bounds for weakly dependent data drawn from sub-exponential distributions, are drawn from white2014asymptotic, merlevede2009bernstein, rio2000inegalites and dedecker2007empirical.
@@
Let $Y_{1},\cdots,Y_{n}$ be a strictly stationary real-valued sequence with common CDF. $F$ and assume the strong mixing coefficients $\upalpha(k)$ in (39) satisfies $\sum_{k\geq 1}\upalpha(k)^{1/2}<\infty$. Define the empirical CDF as per Eq. (35). Then for every $t>0$ and $n\geq 1$
@@
where $C_{\upalpha}=\sum\limits_{k\geq 1}\upalpha(k)^{1/2}<\infty$. In particular, if $\upalpha(k)=0$ for all $k\geq 1$ (the independent case) then $C_{\upalpha}=0$ and (41) reduces to $2e^{-nt^{2}/2}$, which is non-sharp relative to massart1990tight.
@@
In particular, if $\upalpha(k)\equiv 0$ then $C_{\upalpha}=0$ and this reduces to the usual Azuma-Hoeffding bound $\Pr(\lvert\hat{p}_{n}-p\rvert\geq t)\leq 2e^{-nt^{2}/2}$.
@@ B.1 Proof of Lemma 1
Fix $u\in[\varepsilon,1-\varepsilon]$ and $y_{u}:=\hat{F}_{N}^{-1}(u)$. Then $\hat{F}_{N}(y_{u})\geq u$. Hence $F(y_{u})\geq\hat{F}_{N}(y_{u})-\varepsilon\geq u-\varepsilon$, so $F^{-1}(u-\varepsilon)\leq y_{u}$. For any $y<y_{u}$, $\hat{F}_{N}(y)<u$, hence $F(y)\leq\hat{F}_{N}(y)+\varepsilon<u+\varepsilon$, so $F^{-1}(u+\varepsilon)\geq y_{u}$.
B.2 Proof of Lemma 6
From Section 2, Theorem 1 and Remark 1 of dedecker2007empirical, it is known that for any finite measure $\upmu$ and $p\geq 2$:
@@ B.2 Proof of Lemma 6
Next we relate the $\tau$ coefficients to $\upalpha$-mixing. Inequality (4.1) in Section 4.1 of dedecker2007empirical shows:
@@ B.3 Proof of Proposition 1
$\Re\in\bigg[\hat{\delta}_{1}\hat{p}_{1}+L^{(1)}\hat{p}_{0}-U^{(0)}\hat{p}_{1}-\hat{\delta}_{0}\hat{p}_{0},\;\hat{\delta}_{1}\hat{p}_{1}+U^{(1)}\hat{p}_{0}-L^{(0)}\hat{p}_{1}-\hat{\delta}_{0}\hat{p}_{0}\bigg].$ (47)
For $u=m_{0},\cdots,m_{1}$, to simultaneously obtain the $(1-\alpha_{u})\%$ confidence set for both the upper and lower bounds for the identification region (47), we must first find the confidence bands with an appropriate significance level and combine them using Bonferroni inequalities so that the combined confidence set has $100(1-\alpha_{u})\%$ coverage, or
$\Pr\left([l(\Re),u(\Re)]\subseteq[\mathcal{L}(\hat{\mathbf{\theta}}),\mathcal{U}(\hat{\mathbf{\theta}})]\right)\geq 1-\alpha_{u},\quad\text{with}\quad\alpha_{u}=\frac{\alpha}{\overline{\mathcal{M}}},$ (48)
where $\hat{\mathbf{\theta}}=(\hat{\delta}_{1},\hat{\delta}_{0},\hat{p}_{1},\hat{p}_{0})^{\top}$ is a $4\times 1$ vector of estimators and $\mathbb{I}(\Re)=[l(\Re),u(\Re)]$. In other words, we wish to obtain
$\Pr\left(l(\Re)\leq\mathcal{L}(\hat{\theta})\right)\geq 1-\frac{\alpha_{u}}{2},\quad\text{and}\quad\Pr\left(u(\Re)\geq\mathcal{U}(\hat{\theta})\right)\geq 1-\frac{\alpha_{u}}{2},$ (49)
such that
$\Pr\left(l(\Re)\leq\mathcal{L}(\hat{\theta})\cap u(\Re)\geq\mathcal{U}(\hat{\theta})\right)\geq 1-\alpha_{u}.$ (50)
@@
$\begin{split}\Pr\left(l(\Re)\leq\mathcal{L}(\hat{\theta})\cap u(\Re)\geq\mathcal{U}(\hat{\theta})\right)&\geq 1-\Pr\left(l(\Re)>\mathcal{L}(\hat{\theta})\right)-\Pr\left(u(\Re)<\mathcal{U}(\hat{\theta})\right)\\ &\geq 1-\frac{\alpha_{u}}{2}-\frac{\alpha_{u}}{2}\\ &=1-\alpha_{u}.\end{split}$ (51)
Thus,
@@
$\sqrt{N}(\hat{\theta}-\theta)\xrightarrow{d}N(0,\Omega),$ (53)
@@
$\mathcal{L}(\hat{\theta})=\mathcal{L}(\theta)+\nabla\mathcal{L}(\theta)^{\top}(\hat{\theta}-\theta)+o_{p}\left(\lVert\hat{\theta}-\theta\rVert\right),$ (54)
$\mathcal{U}(\hat{\theta})=\mathcal{U}(\theta)+\nabla\mathcal{U}(\theta)^{\top}(\hat{\theta}-\theta)+o_{p}\left(\lVert\hat{\theta}-\theta\rVert\right),$ (55)
@@
$\nabla\mathcal{L}(\theta)=\left(p_{1},\;-p_{0},\;\delta_{1}-U^{(0)},\;L^{(1)}-\delta_{0}\right)^{\top},$ (56)
with $\nabla\mathcal{U}(\theta)$ defined similarly. Since $\hat{\theta}-\theta=O_{p}(N^{-1/2})$, we have $o_{p}(\lVert\hat{\theta}-\theta\rVert)=o_{p}(N^{-1/2})$, and hence
$\mathcal{L}(\hat{\theta})-\mathcal{L}(\theta)=\nabla\mathcal{L}(\theta)^{\top}(\hat{\theta}-\theta)+o_{p}(N^{-1/2}),$ (57)
$\sqrt{N}(\mathcal{L}(\hat{\theta})-\mathcal{L}(\theta))=\nabla\mathcal{L}(\theta)^{\top}\sqrt{N}(\hat{\theta}-\theta)+o_{p}(1).$ (58)
Therefore,
$\sqrt{N}(\mathcal{L}(\hat{\theta})-\mathcal{L}(\theta))\xrightarrow{d}N\left(0,\nabla\mathcal{L}(\theta)^{\top}\Omega\nabla\mathcal{L}(\theta)\right),$ (59)
and similarly
$\sqrt{N}(\mathcal{U}(\hat{\theta})-\mathcal{U}(\theta))\xrightarrow{d}N\left(0,\nabla\mathcal{U}(\theta)^{\top}\Omega\nabla\mathcal{U}(\theta)\right).$ (60)
Consequently,
$\text{Var}\left(\mathcal{L}(\hat{\theta})\right)\approx\frac{1}{N}\nabla\mathcal{L}(\theta)^{\top}\Omega\,\nabla\mathcal{L}(\theta),$ (61)
@@
which determines $\Omega$ in the CLT after appropriate rescaling.
B.4 Proof of Proposition 2
Let the potential outcomes $\{Y_{ijt}^{(k)}\}_{i,j,t}$ be sub-exponential with $\psi_{1}$-norm bounded by $M_{k}$, and assume that for each fixed threshold $\tau_{u}$, the observed outcomes in each arm,
$\{Y_{ijt}:Z_{ijt}(\tau_{u})=k\},\qquad k\in\{0,1\},$
are i.i.d. samples. We wish to show how to obtain the coverage probability
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha$ (63)
for $u=m_{0},\cdots,m_{1}$ and some arbitrary significance level $0<\alpha<1$ when assumption 1 holds. To achieve this, we first need to consider the six “good” events:
$\mathcal{E}_{1}$ $:=\{\lvert\hat{\mu}_{1}-\mu_{1}\rvert\leq t_{1}\},$
$\mathcal{E}_{2}$ $:=\{\lvert\hat{\mu}_{0}-\mu_{0}\rvert\leq t_{2}\},$
$\mathcal{E}_{3}$ $:=\{\lvert\hat{p}_{1}-p_{1}\rvert\leq t_{3}\},$
$\mathcal{E}_{4}$ $:=\{\lvert\hat{p}_{0}-p_{0}\rvert\leq t_{4}\},$
$\mathcal{E}_{5}$ $:=\left\{\sup_{y}\bigg\lvert F_{N_{1}}^{(1)}(y)-F^{(1)}(y)\bigg\rvert\leq t_{5}\right\},$
$\mathcal{E}_{6}$ $:=\left\{\sup_{y}\bigg\lvert F_{N_{0}}^{(0)}(y)-F^{(0)}(y)\bigg\rvert\leq t_{6}\right\}.$
Thus, showing $\Pr\left(\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha_{u}$ for $u=m_{0},\cdots,m_{1}$ is equivalent to showing that the intersection of the events, i.e., $\Pr(\bigcap\limits_{i=1}^{6}\mathcal{E}_{i})\geq 1-\alpha_{u}$. Using De Morgan’s law, it is clear that
$\Pr\left(\bigcap\limits_{i=1}^{6}\mathcal{E}_{i}\right)=1-\Pr\left(\bigcup\limits_{i=1}^{6}\mathcal{E}^{c}_{i}\right),$ (64)
@@ B.4 Proof of Proposition 2
$\geq 1-\sum\limits_{i=1}^{6}\Pr\left(\mathcal{E}_{i}^{c}\right).$ (67)
Hence, showing that the bound (63) holds is equivalent to ensuring that $\sum_{i=1}^{6}\Pr\left(\mathcal{E}_{i}^{c}\right)\leq\alpha_{u}$ for $u=m_{0},\cdots,m_{1}$. The only tools we need are the three inequalities in Lemmas 2-4.
@@
Let $N_{1}$ (resp. $N_{0}$) be the number of observations with $Z(\tau_{u})=1$ (resp. $Z(\tau_{u})=0$). Lemma 2 gives for any $t>0$
@@
The first term is used when $t_{k}\leq M_{k}$ — the “quadratic” regime; otherwise the second, “linear”, term is smaller. With this choice $2\exp[-\log(12/\alpha_{u})]=\alpha_{u}/6$, so $\Pr(\mathcal{E}_{1}^{c})=\Pr(\mathcal{E}_{2}^{c})=\alpha_{u}/6$.
@@
$\Pr\left(\lvert\hat{p}_{k}-p_{k}\rvert\geq t\right)\leq 2\exp[-2N_{k}t^{2}],\qquad k=0,1.$
@@
$\Pr\left(\sup_{y}\lvert F^{(k)}_{N_{k}}(y)-F^{(k)}(y)\rvert>t\right)\leq 2\exp[-2N_{k}t^{2}],\qquad k=0,1.$
@@
On $\cap_{i=1}^{6}\mathcal{E}_{i}$ we have $\lvert\hat{\mu}_{k}-\mu_{k}\rvert\leq t_{k}$, $\lvert\hat{p}_{k}-p_{k}\rvert\leq t_{k+2}$, and $\sup_{y}\lvert F^{(k)}_{N_{k}}(y)-F^{(k)}(y)\rvert\leq t_{k+4}$ for $k=0,1$. By the quantile-sandwich (Lemma 1), for every $p\in[t_{k+4},\,1-t_{k+4}]$,
$F^{-1}(p-t_{k+4})\leq\hat{F}^{-1}_{N_{k}}(p)\leq F^{-1}(p+t_{k+4}).$
Choose any $r_{k}\in\left(0,\frac{1}{2}-t_{k+4}\right]$ and define the data-driven tail endpoints
$L^{(k)}(t_{k+4}):=\hat{F}^{-1}_{N_{k}}(r_{k}+t_{k+4})=Y^{(k)}_{\left(\lceil(r_{k}+t_{k+4})\,N_{k}\rceil\right)},$
and
$U^{(k)}(t_{k+4}):=\hat{F}^{-1}_{N_{k}}(1-r_{k}-t_{k+4})=Y^{(k)}_{\left(\lceil(1-r_{k}-t_{k+4})N_{k}\rceil\right)}.$
Then, on $\mathcal{E}_{k+4}$,
$F^{-1}(r_{k})\leq L^{(k)}(t_{k+4}),\quad U^{(k)}(t_{k+4})\leq F^{-1}(1-r_{k}),$
so $\left[L^{(k)}(t_{k+4}),\;U^{(k)}(t_{k+4})\right]$ are conservative surrogates for the unknown population tail quantiles $\left[F^{(k),-1}(r_{k}),F^{(k),-1}(1-r_{k})\right]$. (When $t_{k+4}\leq\frac{1}{4}$ one may take $r_{k}=t_{k+4}$, which yields the convenient indices $\lceil 2t_{k+4}N_{k}\rceil$ and $\lceil(1-2t_{k+4})N_{k}\rceil$.)
@@
and
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha,$
which establishes (63).
B.5 Proof of Corollary 1
The argument follows Proposition 2 verbatim except that the empirical-CDF events are now one-sided because the lower support is the known constant $\lambda$: